# Contracting with a Mechanism Designer

### v1

@@ Appendix C Missing Proofs
=$\overset{(a)}{=}{}$ $\max\limits_{\alpha\in[0,1]}\sum_{i\in[n]}(1-\alpha)r\cdot\operatorname*{\mathrm{Pr}}\nolimits_{\bm{w}_{\alpha}\sim\bm{F}_{\alpha}}\!\left[{w_{\alpha,i}\geq\theta_{\alpha,i}\wedge\phi_{\alpha,i}(w_{\alpha,i})\geq\max_{j\in[n]}\phi_{\alpha,j}(w_{\alpha,j})}\right]$
=$\overset{(b)}{=}{}$ $\max\limits_{\alpha\in[0,1]}\sum_{i\in[n]}(1-\alpha)r\cdot\operatorname*{\mathrm{Pr}}\nolimits_{\bm{w}\sim\bm{F}}\!\left[{w_{i}\geq(1-\alpha)r+\theta_{\alpha,i}\wedge\phi_{i}(w_{i})\geq\max_{j\in[n]}\phi_{j}(w_{j})}\right]$
=$\overset{(c)}{=}{}$ $\max\limits_{\alpha\in[0,1]}\sum_{i\in[n]}\phi_{i}(z_{\alpha,i})\cdot\operatorname*{\mathrm{Pr}}\nolimits_{\bm{w}\sim\bm{F}}\!\left[{w_{i}\geq z_{\alpha,i}\wedge\phi_{i}(w_{i})\geq\max_{j\in[n]}\phi_{j}(w_{j})}\right]$
=$\overset{(d)}{=}{}$ $\max\limits_{\alpha\in[0,1]}\sum_{i\in[n]}\phi_{i}(z_{\alpha,i})\cdot\operatorname*{\mathrm{Pr}}\nolimits_{\bm{w}\sim\bm{F}}\!\left[{w_{i}\geq z_{\alpha,i}\wedge\phi_{i}(w_{i})=\max_{j\in[n]}\phi_{j}(w_{j})}\right].$
Equality (a) holds since the value distribution $F_{\alpha,i}$ is regular by Lemma C.2 and 2.4. Equality (b) holds because $w_{\alpha,i}=(\alpha-1)r+w_{i}$ and the following fact:
@@
where the first equality holds by definition, the second equality holds because $z_{\alpha,i}=\alpha^{\star}r+\theta_{\alpha,i}$, $F_{\alpha,i}(\theta_{\alpha,i})=F_{i}(z_{\alpha,i})$ and $F^{\prime}_{\alpha,i}(\theta_{\alpha,i})=F^{\prime}_{i}(z_{\alpha,i})$, and the last equality is by the the fact that the reserve price $\theta_{\alpha,i}$ is the zero point of virtual value function $\phi_{\alpha,i}(\cdot)$. Equality (d) holds since by 2.4, agent $i$’s contribution distribution $F_{i}$ is regular. Hence, we have part (ii) of Theorem 4.1.
@@
[…]
=$\overset{(a)}{=}{}$ $\max\limits_{\alpha\in[0,1]}\phi(z_{\alpha})\cdot\sum_{i\in[n]}\operatorname*{\mathrm{Pr}}\nolimits_{\bm{w}\sim F^{\otimes n}}\!\left[{w_{i}\geq\max_{j\in[n]}\left\{z_{\alpha},w_{j}\right\}}\right]$
=$\overset{(b)}{=}{}$ $\max\limits_{\alpha\in[0,1]}\phi(z_{\alpha})(1-F^{n}(z_{\alpha}))$
=$\overset{(c)}{=}{}$ $\max\limits_{z\in\mathtt{supp}(F)}\phi(z)(1-F^{n}(z)),$
@@
Finally, we prove equality (c). By Lemma 4.2, the function $\theta_{\alpha}$ (w.r.t. $\alpha$) is continuous in $[0,1]$. This implies the function $z_{\alpha}$ (w.r.t. $\alpha$) is also continuous in $[0,1]$. Notice that $z_{0}=\theta$ and $z_{1}=r$, where $\theta$ is the monopoly reserve price of contribution distribution $F$. We have $[\theta,r]\subseteq\{z_{\alpha}:\alpha\in[0,1]\}\subseteq\mathtt{supp}(F)$. Hence, equality (c) holds with $\leq$”. On the flip hand, since the contribution distribution $F$ is regular, by the definition of reserve price $\theta$, we have $\phi(z_{\alpha})\leq 0$ if $z_{\alpha}\leq\theta$. Thus, $\phi(z_{\alpha})(1-F^{n}(z_{\alpha}))$ only takes non-negative values if $z_{\alpha}\in[\theta,r]$, leading that equality (c) holds with $\geq$”. Consequently, we have proved equality (c). ∎
C.3 Proof of Lemma 4.2
@@
By the definition of reserve price, we have
Define two functions $\upsilon(\alpha)\triangleq-\theta_{\alpha}+\alpha\bm{\gamma}$ and $\varphi(z)\triangleq z+\frac{G(z)}{G^{\prime}(z)}$. By replacing $\theta_{\alpha}$ in eq. 6 with $-\upsilon(\alpha)+\alpha\bm{\gamma}$, we get that for any $\alpha\in[0,1]$, it holds
Clearly, $\varphi(\upsilon(\cdot))$ is continuous in $[0,1]$. By 2.4, $\varphi(\cdot)$ is continuous in $\mathtt{supp}(G)$. Hence, we have $\upsilon(\cdot)$ is continuous in $[0,1]$. Since $\theta_{\alpha}=-\upsilon(\alpha)+\alpha\bm{\gamma}$, we finally obtain that $\theta_{\alpha}$ is also continuous in $[0,1]$. This completes our proof. ∎
C.4 Proof of Theorem 6.1
See 6.1
@@
We first argue that there exists an optimal contract that is linear. For any optimal contract $\bm{t}^{\star}$, since agents have identical expected reward profiles $\bm{\gamma}$, there exists a linear contract $\alpha^{\star}$ such that

### v2

@@ Appendix C Missing Proofs
$\overset{(a)}{=}{}$ $\max\limits_{\alpha\in[0,1]}\sum_{i\in[n]}(1-\alpha)r\cdot\operatorname*{\mathrm{Pr}}\nolimits_{\bm{w}_{\alpha}\sim\bm{F}_{\alpha}}\!\left[{w_{\alpha,i}\geq\theta_{\alpha,i}\wedge\phi_{\alpha,i}(w_{\alpha,i})\geq\max_{j\in[n]}\phi_{\alpha,j}(w_{\alpha,j})}\right]$
$\overset{(b)}{=}{}$ $\max\limits_{\alpha\in[0,1]}\sum_{i\in[n]}(1-\alpha)r\cdot\operatorname*{\mathrm{Pr}}\nolimits_{\bm{w}\sim\bm{F}}\!\left[{w_{i}\geq(1-\alpha)r+\theta_{\alpha,i}\wedge\phi_{i}(w_{i})\geq\max_{j\in[n]}\phi_{j}(w_{j})}\right]$
$\overset{(c)}{=}{}$ $\max\limits_{\alpha\in[0,1]}\sum_{i\in[n]}\phi_{i}(z_{\alpha,i})\cdot\operatorname*{\mathrm{Pr}}\nolimits_{\bm{w}\sim\bm{F}}\!\left[{w_{i}\geq z_{\alpha,i}\wedge\phi_{i}(w_{i})\geq\max_{j\in[n]}\phi_{j}(w_{j})}\right]$
$\overset{(d)}{=}{}$ $\max\limits_{\alpha\in[0,1]}\sum_{i\in[n]}\phi_{i}(z_{\alpha,i})\cdot\operatorname*{\mathrm{Pr}}\nolimits_{\bm{w}\sim\bm{F}}\!\left[{w_{i}\geq z_{\alpha,i}\wedge\phi_{i}(w_{i})=\max_{j\in[n]}\phi_{j}(w_{j})}\right].$
Equality (a) holds since the value distribution $F_{\alpha,i}$ is regular by Lemma C.4 and 2.3. Equality (b) holds because $w_{\alpha,i}=(\alpha-1)r+w_{i}$ and the following fact:
@@
where the first equality holds by definition, the second equality holds because $z_{\alpha,i}=\alpha^{\star}r+\theta_{\alpha,i}$, $F_{\alpha,i}(\theta_{\alpha,i})=F_{i}(z_{\alpha,i})$ and $F^{\prime}_{\alpha,i}(\theta_{\alpha,i})=F^{\prime}_{i}(z_{\alpha,i})$, and the last equality is by the the fact that the reserve price $\theta_{\alpha,i}$ is the zero point of virtual value function $\phi_{\alpha,i}(\cdot)$. Equality (d) holds since by 2.3, agent $i$’s contribution distribution $F_{i}$ is regular. Hence, we have part (ii) of Theorem 3.1.
@@
$={}$ $\max\limits_{\alpha\in[0,1]}\phi(z_{\alpha})\cdot\sum_{i\in[n]}\operatorname*{\mathrm{Pr}}\nolimits_{\bm{w}\sim F^{\otimes n}}\!\left[{w_{i}\geq z_{\alpha}\wedge\phi(w_{i})=\max_{j\in[n]}\phi(w_{j})}\right]$
$\overset{(a)}{=}{}$ $\max\limits_{\alpha\in[0,1]}\phi(z_{\alpha})\cdot\sum_{i\in[n]}\operatorname*{\mathrm{Pr}}\nolimits_{\bm{w}\sim F^{\otimes n}}\!\left[{w_{i}\geq\max_{j\in[n]}\left\{z_{\alpha},w_{j}\right\}}\right]$
$\overset{(b)}{=}{}$ $\max\limits_{\alpha\in[0,1]}\phi(z_{\alpha})(1-F^{n}(z_{\alpha}))$
$\overset{(c)}{=}{}$ $\max\limits_{z\in\mathtt{supp}(F)}\phi(z)(1-F^{n}(z)),$
@@
Finally, we prove equality (c). By Lemma C.5, the function $\theta_{\alpha}$ (w.r.t. $\alpha$) is continuous in $[0,1]$. This implies the function $z_{\alpha}$ (w.r.t. $\alpha$) is also continuous in $[0,1]$. Notice that $z_{0}=\theta$ and $z_{1}=r$, where $\theta$ is the monopoly reserve price of contribution distribution $F$. We have $\theta,r]\subseteq\{z_{\alpha}:\alpha\in[0,1]\}\subseteq\mathtt{supp}(F)$. Hence, equality (c) holds with $\leq$”. On the flip hand, since the contribution distribution $F$ is regular, by the definition of reserve price $\theta$, we have $\phi(z_{\alpha})\leq 0$ if $z_{\alpha}\leq\theta$. Thus, $\phi(z_{\alpha})(1-F^{n}(z_{\alpha}))$ only takes non-negative values if $z_{\alpha}\in[\theta,r]$, leading that equality (c) holds with $\geq$”. Consequently, we have proved equality (c). ∎
C.3 Proof of Theorem 4.1
In this section, we will prove Theorem 4.1.
See 4.1
Example C.2.
Lemma C.6 (Feng and Jin, 2024, Theorems 1 and 2).
  1. 1.
  2. 2.
Lemma C.7.
Proof.
We first observe that the virtual value function is weakly decreasing, which implies that the monopoly reserve price is weakly increasing. Indeed, if $\phi^{(n)}(z)\geq\phi^{(n+1)}(z)$ holds for any $z\in\mathtt{supp}(F)$, then by the definition of the reserve price, we have
Given that $F^{n}$ is regular, we also notice that $\phi^{(n)}(z)$ is non-decreasing with respect to $z$. This implies that
$\theta^{(n+1)}\geq\theta^{(n)}.$
$\phi^{(n)}(z)=z-\frac{1}{h^{(n)}}=z-\frac{1-F^{n}}{nF^{n-1}F^{\prime}}.$
$\phi^{(n)}(z)\geq\phi^{(n+1)}(z).$
This completes our proof. ∎
Lemma C.8 (Yan, 2011, Theorem 4.1).
To sell a single item among multiple buyers with i.i.d. and regular distributions, the revenue of anonymous pricing is an $\eta$-approximation to that of the Bayesian revenue-optimal mechanism, where $\eta=\frac{{\mathrm{e}}}{{\mathrm{e}}-1}$.
Definition C.3 (The truncated equal-revenue distribution).
One can verify that the monopoly revenue of the truncated equal-revenue distribution $F^{{\dagger}}$ is also $R^{*}$, and $F^{{\dagger}}$ is also regular since the revenue curve of $F^{{\dagger}}$ is concave. Moreover, we notice that the CDF is discontinuous at $\bar{w}$, which corresponds to a point-mass of $R^{*}/\bar{w}$ at value $\bar{w}$.
Lemma C.9.
Proof.
Let $F$ be an arbitrary regular distribution with support $\left[\underline{w},\bar{w}\right]$ and monopoly revenue $R^{*}\in[\underline{w},\bar{w}]$. Observe that the revenue curve of the truncated equal-revenue distribution $F^{{\dagger}}$ is $R^{{\dagger}}(q)=\min\left\{\bar{w}q,R^{*}\right\}$. Consider the revenue curve of distribution $F$, given by $R(q)=zq$, where q=1F(z)q=1-F(z). On the one hand, we know $R(q)\leq\bar{w}q$ since $z\in[\underline{w},\bar{w}]$; on the other hand, we have $R(q)\leq R^{*}$ by the definition of $R^{*}$. It follows that the revenue curve of distribution $F$ is always positioned below that of the truncated equal-revenue distribution $F^{{\dagger}}$, i.e., $R(q)\leq R^{{\dagger}}(q)$ for all $q\in[0,1]$. Consequently, since both $R(\cdot)$ and $R^{{\dagger}}(\cdot)$ are concave, for any $z\in[\underline{w},\bar{w}]$, we have
$1-F(z)\leq 1-F^{{\dagger}}(z).$
Now, we are ready to prove Theorem 4.1.
Proof of Theorem 4.1..
We first show the upper bound parts utilizing the above lemmas, and then show the lower bound part by analyzing Example C.2.
Upper bounding PoDM. We first observe that $F^{n}$ is also regular by Lemma C.6. Thus, we have $\phi^{(n)}(z)\geq\phi^{(n+1)}(z)$ for any $z$ in $\mathtt{supp}(F)$ by Lemma C.7. Then we can obtain that
$U^{*}\geq\max\limits_{q\in[0,1]}q\cdot R^{\prime}(q).$ (4)
Let $K=2\left\lceil\log\kappa\right\rceil$. Divide the quantile space $[0,1]$ of $F^{n}$ into $K+1$ pieces by introducing the following points: $1>q_{0}>q_{1}>\cdots>q_{K}=0$. We set the slope of the revenue curve (the virtual value) $R^{\prime}(q_{0})=0$ and $R^{\prime}(q_{k})=\underline{w}\cdot 2^{k-2\left\lceil\log\kappa\right\rceil+\log\kappa}$ for every index $k\in[K]$. Note that $R^{\prime}(q_{K})=\bar{w}$. One can verify that this construction holds for any regular distribution $F^{n}$ since the virtual value function is non-decreasing. Consequently, for every $q\in[q_{k},q_{k-1}]$ with $k\in[2:K]$, it holds
and for every $q\in[q_{1},q_{0}]$, it holds
Since the support of $F$ is $\underline{w},\bar{w}]$, we know $R(1)\geq\underline{w}$. Thus, we know that monopoly revenue $R^{*}=R(q_{0})\geq R(1)\geq\underline{w}$. Moreover, since $R(q_{0})=0$, we have
$R^{*}$ $={}\int_{0}^{q_{0}}R^{\prime}(q)\cdot{\mathrm{d}}q$
$={}\sum_{k\in[K]}\int_{q_{k}}^{q_{k-1}}R^{\prime}(q)\cdot{\mathrm{d}}q$
where inequality (a) holds by eq. 5 and inequality (b) holds by eq. 6. It follows that
which implies that there exists $k^{{\dagger}}\in[2:K]$ such that
Note that by Lemma C.8, the second-best benchmark satisfies $R^{*}\leq U_{\textsc{sb}}\leq\eta\cdot R^{*}$. Combine eq. 4, we can obtain that
Finally, we consider the case where $1\leq\bar{w}/\underline{w}\leq 4$.
We let $q^{{\dagger}}$ be the largest quantile such that $R^{\prime}(q^{{\dagger}})\geq\underline{w}/2$ and $q^{*}$ be the monopoly quantile such that $R^{*}=R\left(q^{*}\right)$. Notice $q^{{\dagger}}$ is well-defined since $R^{\prime}(0)=\bar{w}$. By definition, we have $0\leq q^{{\dagger}}\leq q^{*}\leq 1$ and $R^{\prime}(q^{{\dagger}})\leq\underline{w}$. Besides, if $R^{\prime}(q^{{\dagger}})\geq\underline{w}/2$, it must hold $q^{{\dagger}}=q^{*}=1$.
Consequently, we can obtain that
Upper bounding PoA. We next prove the upper bound of PoA.
By Lemma C.9, we have that for any monopoly revenue $R^{*}\geq 0$, the largest first-best benchmark for one agent is achieved by the truncated equal-revenue distribution with monopoly revenue equal to $R^{*}$. We can obtain that
$U_{\textsc{fb}}=\int_{0}^{+\infty}\left(1-F^{n}(z)\right)\cdot{\mathrm{d}}z\leq R^{*}\cdot\left(1+\ln\frac{\bar{w}}{R^{*}}\right)\overset{(a)}{\leq}R^{*}\cdot(1+\ln\kappa),$
As a result, we can obtain that for any regular distribution,
Lower bounding PoDM and PoA. We now analyze Example C.2, which shows the lower bound part of the two ratios, i.e., $\mathrm{PoA}=\Omega_{\kappa}\left(\left(\log\kappa\right)^{2}\right)$ and $\mathrm{PoDM}=\Omega_{\kappa}\left(\log\kappa\right)$.
By definition, the virtual value for $z\in[z_{k},z_{k-1})$ is
$\phi(z)=R^{\prime}(q)=2^{K-k+1},$
Next, we consider the PoA. We have
$U_{\textsc{fb}}={}$ $\int_{0}^{z_{0}}1-F(z)\cdot{\mathrm{d}}z$
$\geq{}$ $\sum_{k\in[2:K]}\int_{z_{k}}^{z_{k-1}}1-F(z)\cdot{\mathrm{d}}z$
$\overset{(a)}{=}{}$ $\sum_{k\in[2:K]}\frac{k-2+2^{1-k}}{K}\cdot\ln\left(2+\frac{1}{2^{k-1}-1}\right)$
$\overset{(b)}{\geq}{}$ $\sum_{2\in[K]}\frac{k-2}{K}\cdot\ln 2$
$\overset{(c)}{=}{}$ $\frac{K^{2}-3K+3}{2K}\cdot\ln 2,$
C.4 Proof of Theorem 4.2
@@
Example C.4.
We know that $F_{r}$ is an MHR distribution with $F(\tilde{w})=0$ and F(0)=11/KF(0)=1-1/K.
Lemma C.10 (Aggarwal et al., 2009, Lemma 1).
For any random variable $v$ that follows an MHR distribution $F$ with support in $0,+\infty)$, $1-F(\theta)\geq\frac{1}{{\mathrm{e}}}$, where $\theta$ is the monopoly reserve price of the distribution $F$.
It is important to note that Lemma C.10 holds only when the support of the distribution is non-negative. We next extend this conclusion to cases with arbitrary support sets.
Corollary C.11.
@@
$1-F(\theta)={}$ ${\mathrm{e}}^{-\int_{-\infty}^{\theta}h(z)\cdot{\mathrm{d}}z}$
$={}$ ${\mathrm{e}}^{-\int_{-\infty}^{0}h(z)\cdot{\mathrm{d}}z-\int_{0}^{\theta}h(z)\cdot{\mathrm{d}}z}$
$\overset{(b)}{=}{}$ $\frac{1}{{\mathrm{e}}}(1-F(0))$
$={}$ $\frac{1}{{\mathrm{e}}}\operatorname*{\mathrm{Pr}}\!\left[{v\geq 0}\right].$
A consequence of Corollary C.11 is that the principal’s utility under a linear contract $\alpha$ can be lower-bounded as follows.
Lemma C.12.
@@
where $z=r(1-\alpha)+\theta_{\alpha}$. Notice that by Lemma C.6, the distribution $F^{k}$ is MHR for any $k\in\mathbb{N}$. We denote the monopoly reserve price of $F_{\alpha}^{k}$ as $\theta_{\alpha}^{(k)}$. Based on Corollary C.11, we directly obtain that
$U(\alpha)={}$ $r(1-\alpha)(1-F_{\alpha}^{n}(\theta_{\alpha}))$
$\geq{}$ $r(1-\alpha)(1-F_{\alpha}^{n}(\theta_{\alpha}^{(k)}))$
$\geq{}$ $r(1-\alpha)\left(1-\left(1-\frac{1}{{\mathrm{e}}}\operatorname*{\mathrm{Pr}}\!\left[{v_{\max}^{(k)}\geq 0}\right]\right)^{n/k}\right)$
$={}$ $r(1-\alpha)\left(1-\left(1-\frac{1}{{\mathrm{e}}}\operatorname*{\mathrm{Pr}}\!\left[{w_{\max}^{(k)}\geq r(1-\alpha)}\right]\right)^{n/k}\right)$
$={}$ $r(1-\alpha)\left(1-\left(1-\frac{1}{{\mathrm{e}}}(1-F^{k}(r(1-\alpha)))\right)^{n/k}\right).$
Moreover, when $k=n$, we have
$U(\alpha)\geq\frac{1}{{\mathrm{e}}}r(1-\alpha)(1-F^{n}(r(1-\alpha))).$
And when $k=1$, we have
Lemma C.12 serves as a crucial tool for our subsequent analysis of the two ratios. Specifically, we will repeatedly utilize the expressions for $k=n$ and $k=1$ in our forthcoming proofs.
Lemma C.13 (Dhangwatnotai et al., 2015, Theorem 3.11).
Lemma C.14 (Jin et al., 2019a, Theorem 1).
Proof of Theorem 4.2..
We first show the upper bound parts utilizing the above lemmas and then show the lower bound part by analyzing Example C.4.
$U^{*}\overset{(a)}{=}{}$ $\max_{\alpha\in[0,1]}U(\alpha)$
$\overset{(c)}{=}{}$ $\frac{1}{{\mathrm{e}}}\theta^{(n)}(1-F^{n}(\theta^{(n)}))$
$\overset{(e)}{=}{}$ $\frac{1}{{\mathrm{e}}^{2}}U_{\textsc{fb}},$
where equality (a) holds by definition of $U^{*}$, inequality (b) holds by Lemma C.12, equality (c) holds by definition of $\theta^{(n)}$, inequality (d) holds by Lemma C.13 and equality (e) holds by the definition of $U_{\textsc{fb}}$. Thus we have $\mathrm{PoA}=\frac{U_{\textsc{fb}}}{U^{*}}\leq{\mathrm{e}}^{2}$ by definition.
Upper bounding PoDM. Next, we consider the upper bound of PoDM.
When, $n>1$, by Lemma C.12, we have
It follows that
$\mathrm{PoDM}={}$ $\frac{U_{\textsc{sb}}}{U^{*}}$
$\leq{}$ $\frac{U_{\textsc{sb}}}{\theta^{(n)}\left(1-\left(1-\frac{1}{{\mathrm{e}}}(1-F(\theta^{(n)}))\right)^{n}\right)}$
$={}$ $\frac{U_{\textsc{sb}}}{\theta^{(n)}(1-F^{n}(\theta^{(n)}))}\cdot\frac{1-F^{n}(\theta^{(n)})}{1-\left(1-\frac{1}{{\mathrm{e}}}(1-F(\theta^{(n)}))\right)^{n}}$
$={}$ $\tau_{n}\cdot Z_{n}(F(\theta^{(n)})).$
Note that $Z_{n}(\cdot)$ is non-decreasing in $[0,1]$ and $1-F^{n}(\theta^{(n)})\geq\frac{1}{{\mathrm{e}}}$ by Lemma C.10, we have
Lemma C.14 provides the numerical results of $\tau_{n}$ for different $n$, which allows us to obtain the numerical results of the lower bound of PoDM. We offer the numerical results of the lower bound in Table 3 when $n\leq 44$. When $n\geq 44$, it can be verified that PoDM can never reach $3.033$ since $\tau_{n}$ and $Z_{n}((1-1/{\mathrm{e}})^{1/n}))$ are both decreasing as $n$ increases. Hence the lower bound of $3.033$ is reached when $n=7$.
$n$ 2 3 4 5 6 7 8 9 10 11
12 13 14 15 16 17 18 19 20 21 22
3.032 3.031 3.031 3.031 3.030 3.029 3.027 3.025 3.023 3.022 3.019
23 24 25 26 27 28 29 30 31 32 33
3.017 3.015 3.013 3.011 3.009 3.007 3.005 3.003 3.001 2.998 2.997
34 35 36 37 38 39 40 41 42 43 44
2.995 2.993 2.991 2.989 2.987 2.986 2.984 2.982 2.980 2.979 2.977
Table 3: The upper bounds of PoDM for MHR distributions when $n\leq 44$.
Lower bounding PoDM and PoA. Now we analyze the lower bound part of the two ratios by considering Example C.4. Formally, we show that for every fixed $n$ and any small real $\varepsilon>0$, there exists a parameter $K>1$ and a reward $r$ satisfying that $\mathrm{PoA}\geq{\mathrm{e}}^{2}-\varepsilon$ and $\mathrm{PoDM}\geq{\mathrm{e}}-\varepsilon$.
which does not rely on the parameter $K$.
Observe that for any $z\geq 0$, we have inequality
which gives the lower and upper bounds of $1-F_{r}^{n}(\cdot)$ as
$U^{*}={}$ $\max_{z\in[0,r]}\phi(z)(1-F_{r}^{n}(z))$
$\leq{}$ $\max_{z\in[0,r]}\left(z-1+{\mathrm{e}}^{z-r}\right)\frac{{\mathrm{e}}^{-z}-{\mathrm{e}}^{-r}}{n^{r}}$
$\leq{}$ $\frac{1}{n^{r}}\max_{z\in[0,r]}\left(z-1+{\mathrm{e}}^{z-r}\right){\mathrm{e}}^{-z}$
$={}$ $\frac{1}{n^{r}}({\mathrm{e}}^{-2}+{\mathrm{e}}^{-r}).$
On the other hand,
$U_{\textsc{fb}}={}$ $\int_{0}^{r}(1-F^{n}(z))\cdot{\mathrm{d}}z$
$\geq{}$ $\int_{0}^{r}\left(\frac{{\mathrm{e}}^{-z}-{\mathrm{e}}^{-r}}{n}-\frac{\left({\mathrm{e}}^{-z}-{\mathrm{e}}^{-r}\right)^{2}}{2n^{2}}\right)\cdot{\mathrm{d}}z$
$\geq{}$ $\int_{0}^{r}\left(\frac{{\mathrm{e}}^{-z}-{\mathrm{e}}^{r}}{n}-\frac{{\mathrm{e}}^{-2z}}{2n^{2}}\right)\cdot{\mathrm{d}}z$
$\geq{}$ $\frac{1}{n^{r}}(1-(r+1){\mathrm{e}}^{-r})-\frac{1}{4n^{2r+1}}(1-{\mathrm{e}}^{-2r}),$
and
$U_{\textsc{sb}}={}$ $\int_{0}^{r}\phi(z)(1-F_{r}^{n}(z))\cdot{\mathrm{d}}z$
$\geq{}$ $\int_{0}^{r}\left(z-1+{\mathrm{e}}^{z-r}\right)\left(\frac{{\mathrm{e}}^{-z}-{\mathrm{e}}^{-r}}{n}-\frac{\left({\mathrm{e}}^{-z}-{\mathrm{e}}^{-r}\right)^{2}}{2n^{2}}\right)\cdot{\mathrm{d}}z$
$\geq{}$ $\int_{0}^{r}\left(z-1+{\mathrm{e}}^{z-r}\right)\left(\frac{{\mathrm{e}}^{-z}-{\mathrm{e}}^{-r}}{n}-\frac{{\mathrm{e}}^{-2z}}{2n^{2}}\right)\cdot{\mathrm{d}}z$
$={}$ $\frac{1}{n^{r}}\left({\mathrm{e}}^{-1}-\frac{1}{2}(r^{2}-2r+5){\mathrm{e}}^{-r}+{\mathrm{e}}^{1-2r}\right)-\frac{1}{8n^{2r+1}}\left({\mathrm{e}}^{-2}+{\mathrm{e}}^{-r-1}+(2r+3){\mathrm{e}}^{-2r}\right).$
Therefore, we can finally obtain the lower bounds
and
Finally, for an arbitrary large $r$, we have $\mathrm{PoA}\geq{\mathrm{e}}^{2}-\varepsilon$ and $\mathrm{PoDM}\geq{\mathrm{e}}-\varepsilon$. ∎
C.5 Proof of Theorem 4.3
In this section, we prove Theorem 4.3.
See 4.3
Our analysis relies on the following technical lemmas.
Lemma C.15 (Giannakopoulos et al., 2021, Lemma 3.3).
Lemma C.16.
Proof.
By Lemma C.12, we have:
Observe that the probability distribution function satisfies that
$\bar{F}(z)=\frac{F(z)-F(0)}{1-F(0)}\leq F(z)$
for every $z\in[0,\infty)$. Recall that $\bar{\mu}_{n}$ is the expected value of the first-order statistic derived from $n$ i.i.d. random variables following the distribution $\bar{F}$. We have
Substituting the bounds of $U(\alpha)$ and $U_{\textsc{fb}}$, we obtain that
$={}$ $\delta\left(1-\left(1-\frac{1}{{\mathrm{e}}}(1-F(\delta\cdot\bar{\mu}_{n})\right)^{n}\right)$
$={}$ $\delta\cdot\left(1-\left(1-\frac{1}{{\mathrm{e}}}(1-F(0))(1-\bar{F}(\delta\cdot\bar{\mu}_{n}))\right)^{n}\right).$
Notice that $\bar{F}$ is MHR since $F$ is MHR. According to Lemma C.15, we have
$\bar{F}(\delta\cdot\bar{\mu}_{n})\leq 1-{\mathrm{e}}^{-\delta\cdot\mathcal{H}_{n}}.$
Therefore, we can further obtain that
$\frac{U(\alpha)}{U_{\textsc{fb}}}\geq{}$ $\delta\cdot\left(1-\left(1-\frac{1}{{\mathrm{e}}}(1-F(0)){\mathrm{e}}^{-\delta\cdot\mathcal{H}_{n}}\right)^{n}\right)$
$={}$ $\delta\cdot\left(1-\left(1-(1-F(0)){\mathrm{e}}^{-\delta\mathcal{H}_{n}-1}\right)^{n}\right),$
which completes the proof. ∎
Now we are ready to prove Theorem 4.3.
Proof of Theorem 4.3..
Consider the design of setting $\alpha^{{\dagger}}=1-\frac{\bar{\mu}_{n}}{r}(1-\frac{\ln\mathcal{H}_{n}}{\mathcal{H}_{n}})$. According to the definition of $\delta$ in Lemma C.16, we have $\delta^{{\dagger}}=1-\frac{\ln\mathcal{H}_{n}}{\mathcal{H}_{n}}$. In addition, using the lower bound of $\frac{U(\alpha)}{U_{\textsc{fb}}}$ in Lemma C.16, we obtain that
$\geq{}$ $\left(1-\frac{\ln\mathcal{H}_{n}}{\mathcal{H}_{n}}\right)\left(1-\left(1-\frac{1-F(0)}{{\mathrm{e}}^{2}}\cdot\frac{\ln n}{n}\right)^{n}\right)$
$={}$ $\left(1-\mathcal{O}_{n}\left(\frac{\log\log n}{\log n}\right)\right)\left(1-n^{-\mathcal{O}_{n}\left(1\right)}\right)$
$={}$ $1-\mathcal{O}_{n}\left(\frac{\log\log n}{\log n}\right).$
Finally, since $\mathrm{PoDM}\leq\mathrm{PoA}$, we have
which completes the proof. ∎
C.6 Proof of Theorem 5.1
See 5.1
Proof.
@@
For any optimal contract $\bm{t}^{\star}$, since agents have identical expected reward profiles $\bm{\gamma}$, there exists a linear contract $\alpha^{\star}$ such that